In a special local election, one candidate received 288 votes less than twice the number of votes of the opponent. How many votes did each candidate receive if the total number of votes in the election is 136,395?
Thank you for your help.

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

In a special local election, one candidate received 288 votes less than twice the number of votes of the opponent.
How many votes did each candidate receive if the total number of votes in the election is 136,395?
---
Let one candidate's vote total = "x"
Then the other received 2x-288
-------------------------------------
Equation:
# + # = 136,395
x + 2x - 288 = 136395
3x = 136683
x = 45,561
2x-288 = 2(45561)-288 = 90834
==================================
Cheers,
Stan H.





Answer by mananth(16946)


let the oponent get x votes
the candidate will get 2x-288
x+2x-288=136395
3x=136395+288
x=136683/3
=45561 votes secured by opponent
136395-45561=90834 votes by the candidate